When light is incident in a metal plate electrons are emitted only when the frequency of the light exceeds a certain value. How has this been explained? The maximum kinetic energy of the electrons emitted from a metallic surface is 1.6 x 10^-19J when the frequency of the incident radiation is 7.5 x 10^14Hz. Calculate the minimum frequency of radiation for which electrons will be emitted.
1
Expert's answer
2020-10-20T07:14:16-0400

Minimum frequency of emitted electron will be contributed by difference of maximum KE and incident electron energy

,so energy for minimum frequency = incident electron energy["hf"] - maximum KE

and Energy of incident electron = "hf = h*7.5*10^{14} = 4.97*10^{-19} joule."

So minimum energy available for minimum frequency = "(4.97-1.6)*10^{-19} = 3.37*10^{-19} joule"

Hence minimum frequency = "(3.37*10^{-19})\/h = 0.508*10^{15} = [5.1*10^{14} Hz Answer]"